Find the cumulative distribution function of the random variable X  representing the number of defective in problem a. Then using F(x). Find

        P(X=1) and P (0<X£2)


1
Expert's answer
2021-11-15T11:12:20-0500

The cumulative distribution function of a random variable X is the probability that X will take values less than or equal to specific value x.

It is denoted by "F_X(x)=P(X\le x)"


The probability that X=1 is "\frac{4}{7}"

The probability that X is greater than 0 and less than or equal to 2 is "\frac{5}{7}"